In a school of 750 students,320 are girls,559 students do some kind of sports if 101 girls do no sports how many boys also do no sport
Answer by ankor@dixie-net.com(22740)

In a school of 750 students, 320 are girls, therefore
750 - 320 = 430 are boys
559 students do some kind of sports, therefore
750 - 559 = 191 do no sports
if 101 girls do no sports, how many boys also do no sports
191 - 101 = 90 boys do no sports